OpenStudy (anonymous):

Solve the linear differential equation: (dy/dx) + 2y = e^(2x)

OpenStudy (lgbasallote):

hmm dont you just solve the integrating factor and then do that thingy like \[y(I.F.) = \int Q(x)(I.F.) dx\] thingy

OpenStudy (lgbasallote):

you are familiar with that method right?

OpenStudy (anonymous):

yes, i just can't get the exact answer. do you remember how to do it?

OpenStudy (lgbasallote):

\[I.F. = e^{\int 2dx} \implies e^{2x}\] that's a start

OpenStudy (lgbasallote):

so then \[y(I.F.) = \int Q(x) (I.F.) dx\] ]\[\implies y(e^{2x}) = \int e^{2x} (e^{2x})dx\] \[\implies y(e^{2x}) = \int e^{4x}dx\] that should be straightforward now

OpenStudy (anonymous):

Okay, so there's the form: 1/u(x) \[1\div u\left( x \right) \int\limits_{?}^{?}q( x) u(x)\]

OpenStudy (anonymous):

so you got q(x)u(x)

OpenStudy (anonymous):

or, as you put it, (I.F.)

OpenStudy (lgbasallote):

i guess i did

OpenStudy (anonymous):

so then under that integral, i get \[e^(4x) \div4\]

OpenStudy (lgbasallote):

right \[\Large \frac{e^{4x}}{4}\]

OpenStudy (anonymous):

and then that gets divided by (I.F.), so \[e^4x / e^2x \]

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

dont forget the 4

OpenStudy (lgbasallote):

\[\huge y = \frac{e^{4x}}{4e^{2x}}\] simplify it to get the final answer

OpenStudy (anonymous):

ah okay, the textbook answer simplified the e^4x and e^2x. I got it now. Thanks for walking me through

OpenStudy (lgbasallote):

welcome ^_^